# Hugo Malard1  Michel Olvera1  Stéphane Lathuiliere1  Slim Essid1 1LTCI, Télécom Paris, Institut Polytechnique de Paris

### v1

@@ 3.3 Deep Non-negative Matrix co-Factorization
Soft co-NMF for token decomposition
We assume access to an input video with a corresponding audio signal. For simplicity, let us initially consider the central frame of the video; we will later extend our approach to handle the entire video, incorporating the temporal dimension. Additionally, we assume the availability of trained audio and image backbones. Feeding the audio signal into the audio backbone produces a matrix $X_{A}\in\mathbb{R}^{N_{T}\times C_{A}}_{+}$, where $N_{T}$ is the number of tokens, and $C_{A}$ is the token dimension. Similarly, the image encoder maps an image to a feature representation $X_{I}\in\mathbb{R}^{HW\times C_{I}}_{+}$, where $H$ and $W$ represent the spatial dimensions, and $C_{I}$ is the number of channels. Note that the spatial dimensions are flattened, resulting in a 2D matrix compatible with NMF inputs. Note that our approach diverges from prior works as the image activations encapsulate spatial, rather than temporal, information.
By applying soft co-NMF, we aim at factorizing $X_{A}$ into matrices $U_{A}\in\mathbb{R}^{N_{T}\times K}_{+}$ and $V_{A}\in\mathbb{R}^{K\times C_{A}}_{+}$, along with $X_{I}$ into $U_{I}\in\mathbb{R}^{HW\times K}_{+}$ and $V_{I}\in\mathbb{R}^{K\times C_{I}}_{+}$, where $K$ denotes the predefined number of factors.
To enforce the non-negativity constraints in Eq. (3), we re-parameterize the problem using the sigmoid function $\sigma$ and optimize matrices $\tilde{U}_{A}\in\mathbb{R}^{N_{T}\times K}$ and $\tilde{U}_{I}\in\mathbb{R}^{HW\times K}$, such that $U_{A}=\sigma(\tilde{U}_{A})$ and $U_{I}=\sigma(\tilde{U}_{I})$ with $U_{I}\in[0,1]^{HW\times K}$ and $U_{A}\in[0,1]^{N_{T}\times K}$. Restricting the decomposition to the interval $[0,1]$ rather than $\mathbb{R}_{+}$ offers the advantage of an interpretable, well-calibrated scale. Specifically, the $k$-th column of $U_{I}$ and $U_{A}$ can be interpreted as the activation of the $k^{th}$ factor for each token. In this way, the Hadamard product between the original tokens and a column of $U_{I}$ or $U_{A}$ can be viewed as soft masking, highlighting the portion of the original input that activates the $k^{th}$ factor. Overall, those shifts lead us to challenge the existing soft co-NMF framework, proposing instead that audio and image factors should align, while activations may differ. Thus, we define a specific penalty function to impose semantic closeness between the factor matrices. As the modality-specific factors lie in different spaces, we introduce a new manner to ensure their semantic closeness.
Penalty function
Applying soft co-NMF to a specific problem requires careful definition of the penalty term $P$ in Eq. (3), as it impacts both the quality of the factorization and its interpretability. To achieve a semantically consistent factorization of audio-visual observations, $P$ must effectively capture the semantic similarity between $V_{I}$ and $V_{A}$. However, since the backbones (e.g., CLIP and CLAP) are not aligned, audio and image factors are encoded in two different subspaces, making $L_{1}$ or $L_{2}$ norms unsuitable as penalty functions. Moreover, audio and image representations are not necessarily commensurable: they often differ in dimensionality, precluding direct use of standard distances.
To address this, we propose projecting the audio and visual feature representations into a unified feature space to enable direct comparison. Specifically, we estimate aligned semantic descriptors $\mathcal{D}_{I^{k}}$ and $\mathcal{D}_{A^{k}}$ for the $k^{th}$ column of $U_{I}$ and $U_{A}$, respectively. These descriptors are computed using $J$ semantic anchors, $\{b_{I}^{j}\}_{j=1}^{J}$ and $\{b_{A}^{j}\}_{j=1}^{J}$, which lie in the same spaces as $X_{I}$ and $X_{A}$. Each anchor pair $(b_{I}^{j},b_{A}^{j})$ represents a shared semantic concept, enabling the estimation of aligned semantic descriptors.
Specifically, to estimate aligned semantic anchors, we utilize CLIP [10] and CLAP [32] as image and audio backbones. Since these encoders are aligned with text encoders, we can exploit their respective text embeddings to align identical words across audio and image spaces. Thus, we define $b_{I}^{j}=\mathcal{E}_{I}(t_{j})$ and $b_{A}^{j}=\mathcal{E}_{A}(t_{j})$, where $\mathcal{E}_{I}$ and $\mathcal{E}_{A}$ are the text encoders of CLIP and CLAP, respectively, and $\{t_{j}\}_{j=1}^{J}$ represents a predefined word bank.
@@
Figure 3: CLIP and CLAP encode audio and images in different spaces, complicating distance computation. Our method makes use of semantic anchors to project semantic components (pre-trained representations soft masked by the activations $U_{I}^{k}$ and $U_{A}^{k}$ obtained during the co-factorization) in an audio-visual semantic space where standard distances can be applied. Those distances are used as a penalty function during the decomposition to ensure that the audio and image concepts are the same.
We now describe the process of estimating the semantic descriptors from the matrices $U_{A}$ and $U_{I}$ using the aligned semantic anchors. For simplicity, we detail only the estimation of semantic image descriptors; analogous steps apply to the audio descriptors.
@@
First, we introduce semantic components $\mathcal{C}_{I}^{k}$ associated to each column $k$ of $U_{I}$. Since $U_{I}^{k}$ represents the spatial activation of the $k^{th}$ factor (corresponding to $V_{I}^{k}$), it can be interpreted as a soft segmentation mask for the input $X_{I}$. Applying these soft-masks to $X_{I}$ and averaging the resulting representations extracts the concept associated with the $k^{th}$ factor from the image features. Formally, the $k^{th}$ image semantic component is computed as follows:
@@
$\mathcal{D}_{I^{k}}=\begin{pmatrix}\cos(\mathcal{C}_{I}^{k},b_{I}^{1})\\ \vdots\\ \cos(\mathcal{C}_{I}^{k},b_{I}^{J})\end{pmatrix},$ (5)
where $\cos(.)$ denotes cosine similarity. Intuitively, a descriptor quantifies the alignment of the semantic component with each semantic anchor (illustrated in Figure 3).
Note that $V_{I}^{k}$ could have been used to play the role of semantic components instead of $\mathcal{C}_{I}^{k}$. However, $V_{I}^{k}$ is less resilient to variations of the semantic anchors. Indeed, during the NMF optimization, $V_{I}^{k}$ can take any range of values, potentially degenerating to values similar to those of $b_{I}$, which would be equivalent to performing a simple classification. In contrast, each $\mathcal{C}_{I}^{k}$ is inherently constrained to be a weighted average of elements from the original representations, thus discouraging it from merely replicating values from $b_{I}$. We ablate this choice in Section 4.
Audio semantic descriptors $\mathcal{D}_{A^{k}}$ are computed in the same way as their image counterparts. The penalty function is then defined as the cross-entropy between the image and audio semantic descriptors. However, enforcing similarity across all semantic descriptors may be detrimental, as certain semantic components may be visually present but inaudible, or vice versa. To address this, alignment is applied only to the closest audio and image semantic components. This choice is ablated in Section 4. Using the $L_{2}$ norm for reconstruction, we obtain $V_{A}$, $V_{I}$, $U_{A}$, and $U_{I}$ by solving the following optimization problem:
$\min_{V_{A},V_{I},U_{A},U_{I}\geq 0}\big{[}\beta_{1}\|X_{A}-U_{A}V_{A}\|_{2}^{2}$ (6)
$+\beta_{2}\|X_{I}-U_{I}V_{I}\|_{2}^{2}$
$+\beta_{p}\min_{k}\,\text{CE}(\mathcal{D}_{I^{k}},\mathcal{D}_{A^{k}})\big{]},$
where CE denotes the cross-entropy function.
@@
$\mathcal{R}_{t}=-\beta_{t}\sum_{t=1}^{T}\cos(V_{I_{t}}^{k^{\star}},V_{I_{t+1}}^{k^{\star}})+\cos(V_{A_{t}}^{k^{\star}},V_{A_{t+1}}^{k^{\star}}),$ (7)
where $\beta_{t}\geq 0$ controls the temporal regularization.
@@ 3.4 Improved Segmentation with an Open Vocabulary Segmenter
Our soft co-NMF approach generates a segmentation mask for objects producing sound in the audio input through the $k^{\star}$-th row of $U_{I}$. However, the segmentation accuracy is limited, as the CLIP features in the visual input $X_{I}$ were trained only for image-level semantic alignment, not for precise segmentation. To enhance segmentation quality, we propose to integrate a pre-trained open-vocabulary segmentation model. To maintain compatibility with our NMF framework and keep the approach zero-shot, we select a model based on CLIP that preserves alignment with the text encoder without additional fine-tuning of the CLIP encoder. Here, we use the FC-CLIP model [47], which segments images using CLIP’s image encoder via text prompts encoded using the CLIP text encoder.
A valuable feature of FC-CLIP is its shared CLIP embedding space for both image and text inputs, which allows it to be prompted by vectors from the image space. Leveraging this, we prompt the FC-CLIP decoder with factors identified by our NMF framework. More precisely, in our soft co-NMF decomposition, the rows of matrix $V_{I}$ represent the factors in the image space. Therefore, we can use them to prompt FC-CLIP, treating each factor in $V_{I}$ as a segmentation class. Specifically, $V_{I}^{k^{\star}}$ describes the shared “sounding” factor between audio and visual inputs.
Since $V_{I}^{k^{\star}}$ resides in the CLIP space, it can also be compared to class name embeddings, allowing us to infer the segmented object’s class names and achieve sound prompted image semantic segmentation. The complete pipeline is depicted in Figure 2. Note that, incorporating FC-CLIP is computationally efficient, as the CLIP representation is computed only once before the soft co-NMF decomposition, with only the lightweight FC-CLIP decoder adding minor additional processing overhead (there is $\sim$200M parameters for the CLIP encoder and $\sim$20M for the FC-CLIP decoder).
4 Experimental protocol
Implementation details. To extract the modality-specific representations we utilize the MSCLAP 2023 [10] ($\sim$80M parameters) audio backbone and the OpenCLIP [5] ConvNeXt-Large CLIP trained on LAION 2B [33] ($\sim$200M parameters). For all experiments, we fix the decomposition coefficients to $\beta_{I}=\beta_{2}=1,\beta_{p}=125$, and $K=8$. For the temporal consistency constraint, we set the image and audio frame rate to one and use $\beta_{t}=1$. However, since temporal consistency is specifically designed for single-source segmentation, we set $\beta_{t}=0$ for the multi-source task. Those hyperparameters were selected based on preliminary experiments conducted on the AVSBench validation set. The optimization problem is solved using gradient descent in 1800 steps with a learning rate of 0.25. Importantly, the optimization process induces minimal computational overhead compared to the use of all pre-trained modules (details provided in the supplementary materials).
Datasets. We evaluate our method on five datasets: three for sound-prompted segmentation and two for sound-prompted semantic segmentation. For sound-prompted segmentation, we use: AVS-Bench [49], with binary segmentation maps for audio-visually related pixels, divided into single-source (S4,  5K samples) and multi-source (MS3,  400 samples) datasets. Additionally, we use ADE20k sound-prompted [17], which pairs ADE20k segmentation masks and images[48] with audio from VGGSound [2]. For sound-prompted semantic segmentation, we created ADE Sound Prompted Semantic: a semantic version of the ADE20k sound-prompted dataset, by assigning class labels to segmentation masks (detailed in supp. material). Moreover, we conduct experiments on AVSS [50], ($\sim$11K samples) comprising single and multi-source sounding objects.
Evaluation metrics. We adopt standard segmentation metrics for evaluation: mean Intersection over Union (mIoU) and F-score for the semantic segmentation tasks, and mean Average Precision (mAP) as well as mask Intersection over Union (mask-IoU) and F-score for the binary segmentation tasks (we distinguish between mask and mean IoU to ensure comparability with previous works). Each experiment involving NMF decomposition was conducted three times, with the results reported as the average and standard deviation of the scores.
@@ 5 Results and Discussion
In this section, we compare our approach with other unsupervised methods tackling the task of sound-prompted segmentation. Most of the methods rely on fine-tuning (even if done unsupervisedly) which significantly modifies the original representations of pre-trained models. In contrast, our method is training-free and fully preserves the large generalization ability of the models it builds upon.
For semantic segmentation, we use the class names provided with the dataset as the word bank. For non-semantic segmentation, no predefined class set is needed. We consider two variants of our method: in-domain segmentation, which uses a dataset-specific class set, and out-of-domain segmentation, which relies on a general set of 527 AudioSet[11] tags. For fair comparison, we report results using the general class list, unless otherwise noted.
ADE SP S4 MS3 AVSS ADE SP S.
A-FC-CLIP 45.31 51.52 33.67 11.95 24.95
A+-FC-CLIP 41.89 51.63 31.67 11.82 23.73
Baseline. In the absence of prior work on unsupervised sound-prompted semantic segmentation, and to have a fair competitor that relies on the same pre-trained model, we consider a straightforward yet effective baseline approach: Audio-prompted-FC-CLIP (A-FC-CLIP). This method takes an audio input and its corresponding image, utilizing CLAP to determine the most likely class from the dataset’s class pool. FC-CLIP is then prompted to segment this class within the image. Note that for the baseline we use the dataset-specific class list as we use CLAP as a classifier that should rely on predefined class names.
@@
Training-free mask-IoU $\uparrow$ F-Score $\uparrow$ mask-IoU $\uparrow$ F-Score $\uparrow$
@@
ACL-SSL [31] 59.76 69.03 41.08 46.67
A-FC-CLIP 51.52 57.25 33.67 35.87
TACO 64.04$\pm$ 0.25 71.50$\pm$0.20 50.46 $\pm$3.12 58.11$\pm$2.95
@@
Preliminary experiments: impact of clamping. To test the resilience of CLIP representations to clamping negative values, we experiment with FC-CLIP prompted only with the nonnegative part (via ReLU) of the CLIP text encoding, which is referred to as A+-FC-CLIP. Table 1 shows the performance degradation across all datasets. Except for the ADE SP dataset ($\sim$3.5 point drop), the decrease never exceeds 2 points. This experiment validates the use of clamping, enabling the application of NMF-based methods so as to benefit from their interpretability.
@@ 5.1 Sound-prompted Segmentation
We now discuss experiments on (non-semantic) Sound-Prompted Segmentation. Table 2 compares our approach with other unsupervised methods on the AVS-Bench, in both single and multi-source setups. TACO demonstrates strong performance relative to unsupervised training-based alternatives, outperforming the best competitor method by $\sim$5 and $\sim$9 points of mIoU in S4 and MS3 tasks. Our method also largely outperforms the A-FC-CLIP baseline that uses the same pre-trained models, underlining the effectiveness of our decomposition approach.
In the multi-source setup, while baseline performance drops substantially due to its limitation in handling only a single concept, our method shows robust performance. We hypothesize that robustness arises because $V_{I}^{k^{\star}}$ encodes information about all the sounding objects rather than just one, within the sounding factor The hypothesis is explored qualitatively in Section 5.3.
Table 3 presents performance results on the recently proposed ADE Sound Prompted (ADE SP) dataset, along with variations of our approach on the same dataset to understand its gains. We compare existing methods with TACO variants, and our decomposition without FC-CLIP. Our experiments show that using $U_{I}^{k^{\star}}$ alone from the decomposition leads to better segmentation of the image segments compared to previous audio-visual models. This is noteworthy since all competitors (like DenseAV [17], CAVMAE [14], DAVENet [18], and ImageBind [13]) are fine-tuned for this specific task. In contrast, we use frozen features, decompose them to identify matching factors, and demonstrate that these features contain the spatial information necessary for localization. Predictably, the use of FC-CLIP segmenter substantially refines the segmentation and the scores. Now, we examine the different variants of our approach to understand the large performance gap between TACO and existing methods. In the TACO w/o component variant, using $V_{I}^{k}$ and $V_{A}^{k}$ instead of the soft mask representations as semantic components during the optimization process (i.e., using $\mathcal{D}_{I}^{k}\!=\!V_{I}^{k}$ and $\mathcal{D}_{A}^{k}\!=\!V_{A}^{k}$) leads to degraded performance. In TACO No-min, we force all the concepts to be close, instead of only matching the closest, by using a penalty term where the min operator in (6) is replaced by an average. This results in a noticeable drop in performance. The last variant, TACO w/o penalty, shows the results using $\beta_{p}\!=\!0$ (i.e., without penalization), which is equivalent to performing two independent decompositions and choosing $k^{\star}\!=\!\operatorname{argmin}_{k}CE(\mathcal{D}_{I}^{k},\mathcal{D}_{A}^{k})$. As this decomposition is uni-modal, performance significantly drops.
@@
Word Bank S4 MS3 ADE SP
$U_{i}^{k^{\star}}$ General 27.78$\pm$0.06 28.20$\pm$0.61 32.42$\pm$0.74
Specific 28.76$\pm$0.06 24.94$\pm$0.62 27.74$\pm$0.48
TACO Specific 64.73$\pm$0.16 47.23$\pm$2.25 54.05$\pm$1.40
General 64.04$\pm$0.25 50.45$\pm$3.12 51.57$\pm$0.86
Impact of the Word bank. Table 4 presents the mask-IoU scores on the three used sound-prompted segmentation datasets using either a general class list or one specific to the dataset. Using a specific class list overall improves both TACO results (+1 point on S4 and 3.5 points on ADE SP) and the segmentation using $U_{i}^{k^{\star}}$ (+1 point on S4 and 5 points on ADE SP). On the multi-source MS3 dataset, the general class list performs better. This can be explained by the fact that, when multiple sound sources are present, a broader set of concepts can be beneficial, as one may align with the union of the sources (e.g., Family’ could encompass both a woman and a baby making sounds).

### v3

@@ 3.2 Sem Co-NMF: Revisiting soft co-NMF
Figure 3: As CLIP and CLAP encode audio and images in different spaces, our method employs semantic anchors to project semantic components (representations soft-masked by $U_{I}^{k}$ and $U_{A}^{k}$) in an audio-visual semantic space where standard distances can be considered to compute the penalty function.
@@
We now describe the process of estimating the semantic descriptors from the matrices $U_{A}$ and $U_{I}$ using the aligned semantic anchors. For simplicity, we detail only the estimation of semantic image descriptors; analogous steps apply to the audio descriptors. First, we introduce semantic components $\mathcal{C}_{I}^{k}$ associated to each column $k$ of $U_{I}$. Since $U_{I}^{k}$ represents the spatial activation of the $k^{th}$ factor (corresponding to $V_{I}^{k}$), it can be interpreted as a soft segmentation mask for the input $X_{I}$. Applying these soft-masks to $X_{I}$ and averaging the resulting representations extracts the concept associated with the $k^{th}$ factor from the image features. The $k^{th}$ image semantic component is defined as:
@@
$\mathcal{D}_{I^{k}}=\begin{bmatrix}\cos(\mathcal{C}_{I}^{k},b_{I}^{1}),\ \cdots,\ \cos(\mathcal{C}_{I}^{k},b_{I}^{J})\end{bmatrix}^{T}$ (5)
where $\cos(.)$ denotes cosine similarity. Intuitively, a descriptor quantifies the similarity of the semantic component with each semantic anchor (illustrated in Figure 3). Note that $V_{I}^{k}$ could have been used to play the role of semantic components instead of $\mathcal{C}_{I}^{k}$. However, $V_{I}^{k}$ is less resilient to variations of the semantic anchors. Indeed, during the NMF optimization, $V_{I}^{k}$ can take any range of values, potentially degenerating to values very similar to those of $b_{I}$, which would be equivalent to performing a simple classification. In contrast, each $\mathcal{C}_{I}^{k}$ is inherently constrained to be a weighted average of elements from the original representations, thus discouraging it from merely replicating values from $b_{I}$. We ablate this choice in Section 4. Audio semantic descriptors $\mathcal{D}_{A^{k}}$ are computed similarly to their image counterparts. The penalty function is then defined as the cross-entropy between the image and audio semantic descriptors (alternative distances are also explored in Appendix B). However, enforcing similarity across all semantic descriptors may be detrimental, as certain semantic components may be visually present but inaudible, or vice versa. To address this, alignment is applied only to the closest audio and image semantic components. This choice is ablated in Section 4. Using the $L_{2}$ norm for reconstruction, we obtain $V_{A}$, $V_{I}$, $U_{A}$, and $U_{I}$ by solving the following optimization problem:
$\min_{V_{A},V_{I},U_{A},U_{I}\geq 0}\big{[}\|X_{A}-U_{A}V_{A}\|_{2}^{2}+\|X_{I}-U_{I}V_{I}\|_{2}^{2}+\beta_{p}\min_{k}\,\text{CE}(\mathcal{D}_{I^{k}},\mathcal{D}_{A^{k}})\big{]},$ (6)
where CE denotes the cross-entropy function. The pseudo-code of TACO  is given in Appendix C.
@@
$\mathcal{R}_{t}=-\beta_{temp}\sum_{t=1}^{T}\cos(V_{I_{t}}^{k^{\star}},V_{I_{t+1}}^{k^{\star}})+\cos(V_{A_{t}}^{k^{\star}},V_{A_{t+1}}^{k^{\star}}),$ (7)
where $\beta_{temp}\geq 0$ controls the temporal regularization.
@@ 3.3 Improved Segmentation with an Open Vocabulary Segmenter
Our soft co-NMF approach generates a segmentation mask for objects producing sound in the audio input through the $k^{\star}$-th row of $U_{I}$. However, the segmentation accuracy is limited, as the CLIP features in the visual input $X_{I}$ were trained only for image-level semantic alignment, not for precise segmentation. To enhance segmentation quality, we propose to integrate a pre-trained open-vocabulary segmentation model. To maintain compatibility with our NMF framework and keep the approach zero-shot, we select a model based on CLIP that preserves alignment with the text encoder without additional fine-tuning of the CLIP encoder. Here, we use the FC-CLIP model [50], which segments images using CLIP’s image encoder via text prompts encoded using the CLIP text encoder. A valuable feature of FC-CLIP is its shared CLIP embedding space for both image and text inputs, which allows it to be prompted by vectors from the image space. Leveraging this, we prompt the FC-CLIP decoder with factors identified by our NMF framework. More precisely, in our soft co-NMF decomposition, the rows of matrix $V_{I}$ represent the factors in the image space. Therefore, we can use them to prompt FC-CLIP, treating each factor in $V_{I}$ as a segmentation class. Specifically, $V_{I}^{k^{\star}}$ describes the shared “sounding" factor between audio and visual inputs. Since $V_{I}^{k^{\star}}$ resides in the CLIP space, it can also be compared to class name embeddings, allowing us to infer the segmented object’s class names and achieve sound prompted image semantic segmentation. The complete pipeline is depicted in Figure 3. Note that, incorporating FC-CLIP is computationally efficient, as the CLIP representation is computed only once before the soft co-NMF decomposition, with the lightweight FC-CLIP decoder adding minimal overhead (there is $\sim$200M parameters for the CLIP encoder and $\sim$20M for the FC-CLIP decoder).
@@ 4 Results and Discussion
Implementation details. To extract the modality-specific representations we utilize the MSCLAP 2023 [10] ($\sim$80M parameters) audio backbone and the OpenCLIP [5] ConvNeXt-Large CLIP trained on LAION 2B [36] ($\sim$200M parameters) as they exhibit higher performance and generalization than the aligned audio-visual models in audio and image only tasks [10]. We fix the decomposition coefficients to $\beta_{p}=125$, and $K=8$ and we ablate those choices in Appendix B. All the matrices $U_{A}$,$V_{A}$,$U_{I}$ and $V_{I}$ are initialized randomly from Gaussians. We set the image and audio frame rate to 1 and use $\beta_{temp}=1$ for the temporal consistency constraint. Since it targets single-source segmentation, we set $\beta_{temp}=0$ for the multi-source task. Those hyperparameters were selected based on preliminary experiments conducted on the AVSBench validation set. The optimization problem is solved using gradient descent in 1800 steps with a learning rate of 0.25. Note that, this optimization induces minimal computational overhead as detailed in the Appendix D). For semantic segmentation, we use the class names provided with the dataset as the word bank. For non-semantic segmentation, no predefined class set is needed, we therefore used a general class list: the set of 527 AudioSet[12] tags. We ablate this choice in Appendix B.
Datasets and Metrics. We evaluate our method on five datasets: three for sound-prompted segmentation and two for sound-prompted semantic segmentation. For sound-prompted segmentation, we use: AVS-Bench [53], with binary segmentation maps for audio-visually related pixels, divided into single-source (S4,  5K samples) and multi-source (MS3,  400 samples) datasets. We also use ADE20k sound-prompted [18], which pairs ADE20k segmentation masks and images[51] with audio from VGGSound [3]. For sound-prompted semantic segmentation, we created ADE Sound Prompted Semantic: a semantic version of the ADE20k sound-prompted dataset, by assigning class labels to segmentation masks. We also conduct experiments on AVSS [52], ($\sim$11K samples) comprising single and multi-source sounding objects. Regarding metrics, we adopt standard segmentation metrics for evaluation: mean Intersection over Union (mIoU) and F-score for the semantic segmentation tasks, and mean Average Precision (mAP) as well as mask Intersection over Union (mask-IoU) and F-score for the binary segmentation tasks. Each experiment involving NMF decomposition was conducted three times, with the results reported as the average and standard deviation of the scores. Baseline. In the absence of prior work on unsupervised sound-prompted semantic segmentation, and to have a fair competitor that relies on the same models, we consider a straightforward yet effective baseline approach: Audio-prompted-FC-CLIP (A-FC-CLIP). This method takes an audio input and its corresponding image, utilizing CLAP to determine the most likely class from the dataset’s class pool. FC-CLIP is then prompted to segment this class.
@@
Training-free Segmenter mask-IoU $\uparrow$ F-Score $\uparrow$ mask-IoU $\uparrow$ F-Score $\uparrow$
@@
TACO w/o segmenter ($U_{I}^{k^{\star}}$) 29.68$\pm$0.16 41.91$\pm$0.09 25.88$\pm$0.91 30.72$\pm$0.95
ACL-SSL [34] CLIPSeg 59.76 69.03 41.08 46.67
A-FC-CLIP FC-CLIP 51.52 57.25 33.67 35.87
TACO FC-CLIP 64.04$\pm$ 0.25 71.50$\pm$0.20 43.15 $\pm$0.91 47.5$\pm$0.95
@@
Preliminary experiments: impact of clamping. To investigate the resilience of CLIP representations to clamping negative values, using the baseline setup, we prompted FC-CLIP with only the non-negative part of the CLIP text encoding and evaluated its performance. Results in Appendix A confirm the validity of clamping, supporting the use of NMF-based methods to leverage their interpretability.
In the following subsection, we compare our approach with other unsupervised methods tackling the task of sound-prompted segmentation. Most of the methods rely on fine-tuning (even if done unsupervisedly), which significantly modifies the original representations of pre-trained models. In contrast, our method is training-free and fully preserves the generalization ability of the models it builds upon.
@@ 4.1 Sound-prompted Segmentation
We now discuss experiments on (non-semantic) Sound-Prompted Segmentation. Table 1 compares our approach with other unsupervised methods on the AVS-Bench, in both single and multi-source setups. TACO demonstrates strong performance relative to unsupervised training-based alternatives, outperforming the best competitor method by $\sim$5 and $\sim$2 points of mIoU in S4 and MS3 tasks. Our method also largely outperforms the A-FC-CLIP baseline that uses the same pre-trained models, underlining the effectiveness of our decomposition approach. Interestingly, using $U_{I}^{k^{\star}}$ directly as a segmentation map yields results competitive with state-of-the-art methods, despite our model never being exposed to any audiovisual data (only outperformed by MarginNCE [33] in terms of F-score).
In the multi-source setup, while baseline performance drops substantially due to its limitation in handling only a single concept, our method shows robust performance. We hypothesize that robustness arises because $V_{I}^{k^{\star}}$ encodes information about all the sounding objects rather than just one, within the sounding factor. The hypothesis is explored qualitatively in Section 4.3. Table 2 presents performance results on the recently proposed ADE Sound Prompted (ADE SP) dataset, along with variations of our approach on the same dataset to understand its gains. We compare existing methods with TACO variants and our decomposition without FC-CLIP.As in AVSBench, using only $U_{I}^{k^{\star}}$ from our decomposition yields strong segmentation, outperforming prior audio-visual models. Notably, unlike methods such as DenseAV [18], CAVMAE [15], DAVENet [20], and ImageBind [14], which are trained on audio-image pairs, our approach uses only frozen features. By decomposing these features to identify matching factors, we show that they inherently capture the spatial information needed for localization. Predictably, the use of FC-CLIP segmenter substantially refines the segmentation and the scores. Designing a decomposition specifically tailored to be a compatible prompt for FC-CLIP is a core component of our method and a key direction for achieving significantly better results.
To gain further insight, we now examine variants of TACO. In the TACO w/o component variant, using $V_{I}^{k}$ and $V_{A}^{k}$ instead of the soft mask representations as semantic components during the optimization process (i.e., using $\mathcal{D}_{I}^{k}\!=\!V_{I}^{k}$ and $\mathcal{D}_{A}^{k}\!=\!V_{A}^{k}$) leads to degraded performance. In TACO No-min, we replace the min operator in (6) with an average, penalizing all concept distances instead of just the closest. This results in a noticeable drop in performance. The last variant, TACO w/o penalty, shows the results using $\beta_{p}\!=\!0$ (i.e., without penalization), which is equivalent to performing two independent decompositions and choosing $k^{\star}\!=\!\operatorname{argmin}_{k}CE(\mathcal{D}_{I}^{k},\mathcal{D}_{A}^{k})$ ;resulting in a big performance drop.